FOURIER TRANSFORM INFRARED STUDIES OF THE VIBRATIONAL SPECTRA OF HYDROGEN BEARING SILICON-CARBON CLUSTERS IN Ar AT 10 K ^{a}

\underline{DONGSHENG HAN}, C. M. L. RITTBY, AND W. R. M. GRAHAM, Department of Physics, Texas Christ ian University, Fort Worth, TX 76129.

Fourier transform infrared studies of small hydrogen bearing silicon-carbon clusters trapped in solid Ar have resulted in the identification of new species. Among those so far investigated is the linear SiCCH radical produced from the products of the vacuum ultraviolet photolysis of a mixture of SiH_{4} silane, and C_{2}H_{2} acetylene. Two fundamentals of SiCCH have been abserved, the ν_{2}(σ), carbon-carbon stretching mode at 1989.8 cm^{−1} and the ν_{3}(σ), silicon-carbon stretching mode at 636.0 cm^{−1}. Also investigated is SiCH produced by the photolysis of SiH_{4} and CH_{4} methane. The identification of vibrational modes and determination of structures for these species is based on comparison of measured frequencies, relative intensities, and the results of extensive measurements of D and ^{13}C isotopic shifts with the results of ab initio calculations performed in conjunction with this work .

aWork supported by the Welch Foundation